Recent Log Entries

iocse2 restoration failed

Log Number: 3858637Logbook(s): Tags: Body: restorelib 2.2.1 Target restore date was 2020-11-24 09:29:08 The output from the myRestore execution was as follows: iocse2 system@2020-11-24 09:29 2(2) restored seebpm@2020-11-24 09:29 2773(2773) restored SEE_BURT The following are not available in EPICS IOCSE2:do_fill_wire;S IPM1L10.BURT IPM1L11.BURT IPM1L12.BURT IPM1L13.BURT IPM1L14.BURT IPM1L15.BURT IPM1L16.BURT IPM1L17.BURT IPM1L18.BURT IPM3L10.BURT IPM3L11.BURT IPM3L12.BURT IPM3L13.BURT IPM3L14.BURT IPM3L15.BURT IPM3L16.BURT IPM3L17.BURT IPM3L18.BURT IPM5L10.BURT I

iocse4 restored

Log Number: 3858636Logbook(s): Tags: Body: restorelib 2.2.1 Target restore date was 2020-11-25 09:22:49 The output from the myRestore execution was as follows: iocse4 system@2020-11-25 09:22 2(2) restored seebpm@2020-11-25 09:22 2641(2641) restored SEE_BURT 61(61) restored Execute: /usr/csite/certified/bin/ced2Epics -ced ced -A -s -i iocse4 (success) Procedure return status: Success

iocse2 restoration failed

Log Number: 3858635Logbook(s): Tags: Body: restorelib 2.2.1 Target restore date was 2020-11-23 09:29:08 The output from the myRestore execution was as follows: iocse2 system@2020-11-23 09:29 2(2) restored seebpm@2020-11-23 09:29 2773(2773) restored SEE_BURT The following are not available in EPICS IOCSE2:do_fill_wire;S IPM1L10.BURT IPM1L11.BURT IPM1L12.BURT IPM1L13.BURT IPM1L14.BURT IPM1L15.BURT IPM1L16.BURT IPM1L17.BURT IPM1L18.BURT IPM3L10.BURT IPM3L11.BURT IPM3L12.BURT IPM3L13.BURT IPM3L14.BURT IPM3L15.BURT IPM3L16.BURT IPM3L17.BURT IPM3L18.BURT IPM5L10.BURT I

iocse2 restoration failed

Log Number: 3858634Logbook(s): Tags: Body: restorelib 2.2.1 Target restore date was 2020-11-25 09:21:19 The output from the myRestore execution was as follows: iocse2 system@2020-11-25 09:21 2(2) restored seebpm@2020-11-25 09:21 The following are not available in EPICS IOCSE2:CTsub.STR1 IPM1L10.YNIX IPM1L11.YNIX IPM1L12.YNIX IPM1L13.YNIX IPM1L14.YNIX IPM1L15.YNIX IPM1L16.YNIX IPM1L17.YNIX IPM1L18.YNIX IPM3L10.YNIX IPM3L11.YNIX IPM3L12.YNIX IPM3L13.YNIX IPM3L14.YNIX IPM3L15.YNIX IPM3L16.YNIX IPM3L17.YNIX IPM3L18.YNIX IPM5L10.YNIX IPM5L11.YNIX IPM5L12.YNIX IPM5L13.YNIX I

COMMENT ON Injector Upgrade - Phase 1 Installation (Gun to Chopper)

Log Number: 3858630Logbook(s): Body: 

Day #48, Week 10 Day 3, Plans:

Tunnel --

InjSB --
* PSS Upgrades, mostly at IN02{BNI,29,30}, but
occasionally in others throughout building for hardware
installation. (Beizer)
* CIS rack work, IN01B{00-05} (Windham)

Test Lab --
* Documenting UITF 15deg dipole controls for
reproduction at CEBAF (Windham)

Toolbox attended by: Suleiman, Mamun, Croke, Bullard,
Windham

iocnl22rffcc1 restored

Log Number: 3858628Logbook(s): Tags: Body: restorelib 2.2.1 Target restore date was 2020-11-25 08:27:31 The output from the myRestore execution was as follows: iocnl22rffcc1 Rf12Fcc Pause: 60 second(s) Execute: /cs/opshome/burt/REBOOT/Rf12Fcc/DLOAD_ICAL.csh iocnl22rffcc1 (success) c100_rffcc@2020-11-25 08:27 34(34) restored Execute: burtwb -r 3 -f /cs/opshome/burt/REBOOT/Rf12Fcc/snap-ecal/iocnl22rffcc1-ecal.snap (success) Execute: burtwb -r 3 -f /cs/opshome/burt/REBOOT/Rf12Fcc/snap-exp/iocnl22rffcc1-exp.snap (success) Execute: /cs/opshome/burt/REBOOT/Rf12Fcc/GGTmyRestore.csh iocnl22

COMMENT ON Injector Upgrade - Phase 1 Installation (Gun to Chopper)

Log Number: 3858603Logbook(s): Body: 

Day #47, Week 10 Day 2, Wrap-Up:

AIPINJ --
* A1 aperture motor and switch terminations in SB
completed

Other SAD Injector Work --
* PSS upgrades continue --
** Hall D PLCs in InjSB power and I/O cables installed.
3858584
** Relay card re-terminations completed and tested.
3858585

UITF/HDIce day5 run3 summary (Day shift)

Log Number: 3858595Logbook(s): Body: 

Day5 of run3, Tuesday November 24, 2020

A good day, the machine was mostly very reliable and we delivered 250 pA to the polarized HD target for ~ 8 hours uninterrupted throughout Day shift. The program will continue this way through Swing shift.

UITF: field emission 8 hour strip tool

Log Number: 3858586Logbook(s): Entry Makers: poelkerBody: 

8 hour strip tool of the gun field emission, it's constant. Photogun lifetime is fine, attenuator is sitting near 150 to provide HDIce target 250 pA. At such small current, we expect the gun to run forever, but the field emission can influence lifetime, fortunately, it doesn't seem to be a problem for our low current HDIce program

5 hours at 250pA polarized HD target: monitor A5 and A6 interception to look for orbit drifts

Log Number: 3858580Logbook(s): Entry Makers: poelkerBody: 

5 hours of beam at 250 pA to polarized HD target. You can see in the graph below one beam trip at ~ 13:46 which was caused by a quad magnet mismatch. WE restored beam and noticed the beam was not passing through A6 properly, with half the beam clipped. THis means the orbit drifted, and beam got lost on A6, and the attenuator walked up to maintain desired 250 pA current. After putting the beam back on the golden orbit with magnet MHBM904A V, the interception went away and the laser attenuator reduced to the value at the beggining of the run.

COMMENT ON Accelerator Control re-survey

Log Number: 3858576Logbook(s): Body: 

Crew 2: TS, JCr. East Arc dipole re-fid work. MJD1A02 &
1A01 are finished.

MJD1A01 was a full stack (Full stacks are checked for
leaning stands).

Moved back to MJD1A05, will check for leaning.

North linac vacuum activities

Log Number: 3858563Logbook(s): Body: 

1L02. Installed waveguide ceramic warm windows at cavities 1,2,3 & 4.
1L04. Insulating vacuum rough leak test okay.

3+ hours at 250pA, polarized HD target run

Log Number: 3858549Logbook(s): Entry Makers: poelkerBody: 

More than 3 hours into the 250pA run on polarized HD target. FE constant.

Team HDice makes periodic NMR meausurements of polarization

No trips the entire time, things are going well

Pages